A carabao can plough a farmer's land in 6 hours. A tractor can plough the land in 4 hours. The farmer used the carabao to plough the land for 2 hours, after which, another farmer decided to help him using the tractor. How long did it take the farmers tofinish ploughing the land?
Answer by Theo(13342)

r * t = q

r is the rate
t is the time
q is the quantity.

the quantity in this case is one ploughed field.

the carabao can do it in 6 hours, therefore r * t = q becomes r * 6 = 1.
solve for r to get r = 1/6.

the tractor can do it in 4 hours, therefore r * t = q becomes r * 4 = 1.
solve for r to get r = 1/4.

the rate of the carabao is 1/6 of the field in 1 hour.
the rate of the tractor is 1/4 of the field in 1 hour.

the farmer uses the carabao to plough the field for 2 hours.

r * t = q becomes 1/6 * 2 = q.
solve for q to get q = 1/6 * 2 = 1/6 = 1/3.

in 2 hours, the farmer has ploughed 1/3 of the field.

that leave 2/3 of the field to still be ploughed.

another farmer decides to help him using the tractor.

the last 2/3 of the field is then ploughed using the carabao and the tractor.

when both are used, their rates are additive.

the formula of r * t = q becomes (1/6 + 1/4) * t = 2/3

simplify to get 5/12 * t = 2/3

solve for t to get t = 2/3 * 12/5 = 8/5.

the remaining 2/3 of the field is ploughed in 8/5 = 1 and 3/5 hours.

that's your solution.

r * t = q becomes 1/6 * 2 + 5/12 * 8/5 = q
1/6 * 2 is with the carabao only.
5/12 * 8/5 is with both the carabao and the tractor.
solve for q to get q = 1
the whole field is ploughed in 2 hours and 8/5 hours = 3 and 3/5 hours.
that was 2 hours using the carabao and 1 and 3/5 hours using both the tractor and the carabao.

any questions, write to dtheophilis@gmail.com